OpenStudy (jilltalks):

A cone has a volume of 12 cubic inches. What is the volume of a cylinder that the cone fits exactly inside of? 6 in3 24 in3 36 in3 48 in3

OpenStudy (11study11):

I think it is C but I am not 100% sure.

OpenStudy (jilltalks):

Can yo tell me why you think it is C?

OpenStudy (jilltalks):

*you

OpenStudy (11study11):

Because is you divide 36 by 3 you get 12 and it is cubed (3)

OpenStudy (jilltalks):

okay thank you

OpenStudy (legomyego180):

So the trick here is writing out your formulas

OpenStudy (jilltalks):

yes, thank you

OpenStudy (legomyego180):

\[V _{cone}=\frac{ 1 }{ 3 }\pi*r^2*h=12\] \[V _{cyllinder}=\pi*r^2*h\] Do you see where to go from here?